# Cyclic Representations of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ and its Borel Subalgebras at Roots of Unity and Q-operators

### v2

@@ 4.1. Quantum Space $V^{\otimes M}$
(which can be proved either by appealing to uniqueness of the corresponding $U_{q}(\mathfrak{b}_{+})$ intertwiner or just by checking the relation), or directly from Propositions 4.2 and 4.3 $(ii)$ plus the observation that $T_{\varphi}$ commutes with $Q_{\rho_{r}}$ and $Q_{\bar{\rho}_{s}}$. Thus we have that $T_{\Omega_{rs}}(w)$ satisfies the same relations (4.6) as the Q operator of the 6-vertex model. This fact was observed in [BS90].
@@ 4.2. Quantum Space $W^{\otimes M}$
$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$z$${\mathcal{T}}(z)=\hbox{Tr}\left({\bf L}_{\Omega_{ss^{\prime}}}^{M}(z)\cdots{\bf L}_{\Omega_{ss^{\prime}}}^{2}(z){\bf L}^{1}_{\Omega_{ss^{\prime}}}(z)\right)\quad\sim$
@@
${\mathcal{T}}^{CP}=\operatorname*{Tr}_{W^{0}}(R(rr^{\prime};ss^{\prime})^{0M}\cdots R(rr^{\prime};ss^{\prime})^{02}R(rr^{\prime};ss^{\prime})^{01}).$
@@
(4.8) ${\mathcal{T}}^{CP}=\operatorname*{Tr}_{W^{0}}({\mathcal{A}}_{rs;s^{\prime}}\circ P^{(M)}\circ{\mathcal{B}}_{r^{\prime};ss^{\prime}}),$
@@
It was shown in [BS90] that there was a relation between the Q-operator of the $\tau_{2}$ model (found in [BS90] by explicitly solving of TQ relations) and the half-monodromy matrix of the chiral Potts model. This observation and its algebraic origin are reflected and explained in the current paper by three results: the definition ${\mathcal{Q}}_{r^{\prime};ss^{\prime}}=\hbox{Tr}_{W}\left({\mathcal{B}}_{r^{\prime};ss^{\prime}}\right)$; Proposition 4.4; and the identification (4.8).
@@ 5. Discussion
As mentioned in Section 1, another potential use of our results is in the analysis of open systems. Q-operators for some generic-$q$ open systems have been constructed in terms of infinite-dimensional $U_{q}(\mathfrak{b}_{+})$ representations in [FS15, BT18, VW20, Tsu21, CVW24]. The reason this construction is possible is that the recently formulated universal K-matrix for quantum affine algebras [AV22b, AV22a, AV24] is itself intimately related to the Borel algebras, as well as to coideal subalgebras of the underlying quantum affine algebra. Convergence of the Q-operator is a subtle issue for generic-$q$ open systems; in closed systems, convergence can be ensured by inserting regularizing Cartan elements inside the trace, but this is not possible in open systems as integrability is destroyed by such an insertion. Convergence has to demonstrated for open systems by using convergence properties of $q$-special functions in different regions, and TQ relations must be interpreted in this light. A key advantage of dealing with $q^{N}=1$ is that the $U_{q}(\mathfrak{b}_{+})$ cyclic representations we consider are finite dimensional, and so Q-operators are immediately convergent. Another advantage is that it is simpler to solve reflection equations in the finite rather than infinite-dimensional case for the most general boundary conditions. Thus we expect that the results of the current paper as well as higher-rank generalizations will be a useful tool/test-bed in completing the algebraic description of $Q$ operators for open systems.

### v3

@@ 4.1. Quantum Space $V^{\otimes M}$
(which can be proved either by appealing to uniqueness of the corresponding $U_{q}(\mathfrak{b}_{+})$ intertwiner or just by checking the relation), or directly from Propositions 4.2 and 4.3 $(ii)$ plus the observation that $T_{\varphi_{c}}$ commutes with $Q_{\rho_{r}}$ and $Q_{\bar{\rho}_{s}}$. Thus we have that $T_{\Omega_{rs}}(w)$ satisfies the same relations (4.6) as the Q operator of the 6-vertex model. This fact was observed in [BS90].
@@ 4.2. Quantum Space $W^{\otimes M}$
$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$s\,$$s^{\prime}$$z$${\mathcal{T}}(z)=\hbox{Tr}\left((\pi_{z}(t_{1}^{\alpha})\otimes\mathbb{I})\,{\bf L}_{\Omega_{ss^{\prime}}}^{M}(z)\cdots{\bf L}_{\Omega_{ss^{\prime}}}^{2}(z){\bf L}^{1}_{\Omega_{ss^{\prime}}}(z)\right)\quad\sim$
@@
${\mathcal{T}}^{CP}=\operatorname*{Tr}_{W^{0}}((\Omega_{rr^{\prime}}(t_{1}^{\alpha})\otimes\mathbb{I})R(rr^{\prime};ss^{\prime})^{0M}\cdots R(rr^{\prime};ss^{\prime})^{02}R(rr^{\prime};ss^{\prime})^{01}).$
@@
(4.8) ${\mathcal{T}}^{CP}=\operatorname*{Tr}_{W^{0}}((\Omega_{rr^{\prime}}(t_{1}^{\alpha})\otimes\mathbb{I})\,{\mathcal{A}}_{rs;s^{\prime}}\circ P^{(M)}\circ{\mathcal{B}}_{r^{\prime};ss^{\prime}}),$
@@
It was shown in [BS90] that there was a relation between the Q-operator of the $\tau_{2}$ model (found in [BS90] by explicitly solving of TQ relations) and the half-monodromy matrix of the chiral Potts model. This observation and its algebraic origin are reflected and explained in the current paper by three results: the definition ${\mathcal{Q}}_{r^{\prime};ss^{\prime}}=\hbox{Tr}_{W}\left((\Omega_{rr^{\prime}}(t_{1}^{\alpha})\otimes\mathbb{I}){\mathcal{B}}_{r^{\prime};ss^{\prime}}\right)$; Proposition 4.4; and the identification (4.8).
@@ 5. Discussion
As mentioned in Section 1, another potential use of our results is in the analysis of open systems. Q-operators for some generic-$q$ open systems have been constructed in terms of infinite-dimensional $U_{q}(\mathfrak{b}_{+})$ representations in [FS15, BT18, VW20, Tsu21, CVW24]. The reason this construction is possible is that the recently formulated universal K-matrix for quantum affine algebras [AV22b, AV22a, AV24] is itself intimately related to the Borel algebras, as well as to coideal subalgebras of the underlying quantum affine algebra. One barrier to developing this approach and constructing explicit Q-operators for open systems with arbitrary boundary conditions is the complexity of solving infinite-dimensional reflection equations. It is hoped that the finite-dimensional cyclic $U_{q}(\mathfrak{b}_{+})$ representations introduced in this paper will be of use in this context.